# Faster Vision Mamba is Rebuilt in Minutes via Merged Token Re-training

### v1

@@ Abstract
Vision Mamba (e.g., Vim [54]) has successfully been integrated into computer vision, and token reduction has yielded promising outcomes in Vision Transformers (ViTs). However, token reduction performs less effectively on Vision Mamba compared to ViTs. Pruning informative tokens in Mamba leads to a high loss of key knowledge and bad performance. This makes it not a good solution for enhancing efficiency in Mamba. Token merging, which preserves more token information than pruning, has demonstrated commendable performance in ViTs. Nevertheless, vanilla merging performance decreases as the reduction ratio increases either, failing to maintain the key knowledge in Mamba. Re-training the token-reduced model enhances the performance of Mamba, by effectively rebuilding the key knowledge. Empirically, pruned Vims only drop up to 0.9% accuracy on ImageNet-1K, recovered by our proposed framework R-MeeTo in our main evaluation. We show how simple and effective the fast recovery can be achieved at minute-level, in particular, a 35.9% accuracy spike over 3 epochs of training on Vim-Ti. Moreover, Vim-Ti/S/B are re-trained within 5/7/17 minutes, and Vim-S only drop 1.3% with 1.2$\times$ (up to 1.5 $\times$) speed up in inference.
@@ 1 Introduction
Vision Mambas (e.g., Vim [54]) have successfully introduced Mamba [13] into computer vision, achieving promising results [54, 28, 34, 48]. The success of Vim and its follow-up works (Vims) are closely tied to the SSM model’s efficient sequence processing.
@@
Token reduction performances well in input-agnostic Transformer [2]. However, in Mamba, its sequential dependency implies that token reduction varies from Transformer. Therefore, comparative analyses of the performance between Transformer and Mamba are conducted in Fig. 1(a). These present a challenging issue that token pruning operation, designed on ViTs, performs less effectively on Vim.
@@
A closer look at the difference between Transformer and Mamba is taken, and analyses are as follows: after a given time $t$, the token in Mamba contains more general knowledge than that in Transformer, due to the enrichment effect of SSM as Theorem 5. As a result, the tokens at the end of a sequence share the most general knowledge. Pruning these informative tokens leads to the high loss of general knowledge and performance drop. So pruning is not a good solution to make Mamba more efficient.
@@ 2.2 Discussions
The key knowledge (i.e., specific and general knowledge) is reduced by token reduction, and further the remnant tokens and their imbalance lead to performance drop. As shown in Fig. 2, the essential causes are: 1) a large amount of general knowledge is irreparably reduced; 2) specific knowledge keeping ratio is low and imbalanced. Further experiments support our theorem, if we shuffle token after reduction, only Mamba dropping, which means tokens in Transformer is not sensitive to its order of tokens’ indexes. Moreover, token reduction’s disruption to sequential dependency is one of the reasons for performance dropping, due to the knowledge embedded in the tokens’ sequence.
@@ 2.3 Detailed Analyses and Related Theorems
Figure 2: Analysis’ sketch: Mamba is sensitive to token reduction. Experiments about i) token reduction are conducted with DeiT-S [45] (Transformer) and Vim-S [54] (Mamba) on ImageNet-1K [8]. The reduction ratios in the experiment about ii) shuffled tokens are 0.14 for Vim-Ti and 0.31 for Vim-S/Vim-B (see Sec. 3.4 for more details about ablation). Shuffle strategy is odd-even shuffle: [0,1,2,3]→[0,2], [1,3]→[0,2,1,3]. The empirical results of $I(X;Y)$, the mutual information between inputs $X$ and outputs $Y$ of the Attention Block and SSM, are measured by MINE [1] on the middle layers of DeiT-S and Vim-S ($7\text{-th}/12$ layers and the $14\text{-th}/24$ layers respectively.)
@@
Thus, the effect of external factors to performance is ruled out, according to Remark 1, and we therefore study the inherent differences between these modules’ mechanisms.
@@
which can be simply obtained by the commutative of mutual information.
@@ 3.2 Comparative Experiments
Table 2: Comparison between different token reduction methods on the performance of Vim-Ti, Vim-S and Vim-B in ImageNet-1K classification. R-MeeTo (ours) consistently achieves higher top-1 accuracy (%) than competing methods across various scales of Vims and VideoMs while maintaining comparable FLOPs.
@@ 5 Conclusion
Contribution Statement
In this project, mingjia engaged in most of the presentation, figures and tables, project instruction, theoretical derivation, and participated in code review and tuning. Yuhao mainly engaged in experimental tuning, analyses, reproduction, open source, and experimental pictures. Ruiji finished the original code and engaged in part of the experimental design, tuning and information collection.

### v4

@@ Abstract
Vision Mamba has shown close to state of the art performance on computer vision tasks, drawing much interest in increasing it’s efficiency. A promising approach is token reduction (that has been successfully implemented in ViTs). Pruning informative tokens in Mamba leads to a high loss of key knowledge and degraded performance. An alternative, of merging tokens preserves more information than pruning, also suffers for large compression ratios. Our key insight is that a quick round of retraining after token merging yeilds robust results across various compression ratios Empirically, pruned Vims only drop up to 0.9% accuracy on ImageNet-1K, recovered by our proposed framework R-MeeTo in our main evaluation. We show how simple and effective the fast recovery can be achieved at minute-level, in particular, a 35.9% accuracy spike over 3 epochs of training on Vim-Ti. Moreover, Vim-Ti/S/B are re-trained within 5/7/17 minutes, and Vim-S only drops 1.3% with 1.2$\times$ (up to 1.5 $\times$) speed up in inference.
@@ 1 Introduction
Vision Mambas (e.g., Vim [55]) have successfully introduced Mamba [13] into computer vision, achieving promising results [55, 28, 34, 48]. The success of Vim and its follow-up works (Vims) are closely tied to the SSM [13] model’s efficient sequence processing.
@@
Token reduction performances well in input-agnostic Transformer [2], while that on Mamba is unexplored. Therefore, comparative analyses of the performance between Transformer and Mamba are conducted in Fig. 1(a). These present a challenging issue that token pruning operation, designed on ViTs, performs less effectively on Vim.
@@
Mamba’s sequential dependency implies that token reduction varies from Transformer. A closer look at the difference between Transformer and Mamba: after a given time $t$, the token in Mamba contains more general knowledge than that in Transformer, due to the enrichment effect of SSM: tokens are asymmetric in the amount of information they keep (Theorem 5). As a result, the tokens at the end of a sequence share the most general knowledge. Pruning these informative tokens leads to the high loss of general knowledge and performance drop. So pruning is not a good solution to make Mamba more efficient.
@@ 2.2 Discussions
The key knowledge (i.e., specific and general knowledge) is reduced by token reduction, and further the remnant tokens and their imbalance lead to performance drop. As shown in Fig. 2, the essential causes are: 1) a large amount of general knowledge is irreparably reduced; 2) specific knowledge keeping ratio is low and imbalanced. Further experiments in Fig. 3 support our theorem, if we shuffle token after reduction, only Mamba dropping, which means tokens in Transformer is not sensitive to its order of tokens’ indexes. Moreover, token reduction’s disruption to sequential dependency is one of the reasons for performance dropping, due to the knowledge embedded in the tokens’ sequence.
@@ 2.3 Detailed Analyses and Related Theorems
Figure 2: Analysis’ sketch: Mamba is sensitive to token reduction.
Figure 3: Supporting facts. 1) The empirical results of $I(X;Y)$, the mutual information between inputs $X$ and outputs $Y$. Mamba is sensitive to token order. 2) Only Mamba’s performance drops if we further Shuffle Tokens before re-training. The Attention Block and SSM are measured by MINE [1] on the middle layers of DeiT-S and Vim-S ($7\text{-th}/12$ layers and the $14\text{-th}/24$ layers respectively.) Experiments about i) token reduction are conducted with DeiT-S [45] (Transformer) and Vim-S [55] (Mamba) on ImageNet-1K [8]. The reduction ratios in the experiment about ii) shuffled tokens are 0.14 for Vim-Ti and 0.31 for Vim-S/Vim-B (see Sec. 3.4 for more details about ablation). Shuffle strategy is odd-even shuffle: [0,1,2,3]→[0,2], [1,3]→[0,2,1,3].
@@
Thus, the effect of external factors to performance is ruled out with given assumptions, according to Remark 1, and we therefore study the inherent effects and differences between these modules’ mechanisms as followings.
@@
which can be simply obtained by the same derivation as forward SSM and the commutative of mutual information.
@@ 3.2 Comparative Experiments
PlainM-L1 PlainM-L2 PlainM-L3 PlainM-L1 PlainM-L2 PlainM-L3
PlainM (baseline) [49] 77.9 0.0 81.6 0.0 82.3 0.00 3.00 0.00 8.10 0.00 14.4 0.0
Token Recognition [26] 75.0 2.9$\downarrow$ 78.3 2.7$\downarrow$ 78.9 3.4$\downarrow$ 2.44 0.56$\downarrow$ 6.22 1.88$\downarrow$ 8.35 6.05$\downarrow$
Hidden State Alignment [52] 77.4 0.5$\downarrow$ 81.0 0.6$\downarrow$ 81.7 0.6$\downarrow$ 2.46 0.54$\downarrow$ 6.27 1.83$\downarrow$ 8.44 5.96$\downarrow$
R-MeeTo (ours) 77.3 0.6$\downarrow$ 81.4 0.2$\downarrow$ 82.1 0.2$\downarrow$ 2.46 0.54$\downarrow$ 6.29 1.81$\downarrow$ 8.46 5.94$\downarrow$
@@
Table 2: Comparison between different token reduction methods on the performance of Vim-Ti, Vim-S and Vim-B in ImageNet-1K classification. R-MeeTo (ours) consistently achieves higher top-1 accuracy (%) than competing methods across various scales of Vims and VideoMs while maintaining comparable FLOPs. Experiments about PlainM, i.e., PlainMamba, are conducted on resources from [48].